CH/Agen/93/06

Agenda

Charissa Group Meeting

To be held on: Wednesday, 17 Nov 1993, 11.30 am To be held at: Birmingham

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

Agenda

Meeting to Agree Grant Submission

1. Minutes of Meeting of 9 September 1993
2. Matters Arising / Actions
3. Debriefing on ANU running
4. Discussion of Grant Submission
5. Any other business
6. Date of next meeting

W.N. Catford (Secretary) 16 November 1993